# Specific heat study

### v1

@@ III.1 Ferromagnets
Figure 2: Specific heat $c(T)$ for the $S=1/2$ Heisenberg ferromagnet on four three-dimensional lattices. Quantum Monte Carlo simulations for the number of unit cells ${\cal L}=40$ (simple cubic) and ${\cal L}=30$ (pyrochlore, diamond, hyperkagome) and a small symmetry breaking field $h=10^{-4}$. Inset presents $c$ as a function of $T/T_{\rm C}$.
Our findings for ferromagnets are collected in the middle column of Table 1 and in Fig. 2. As is generally known, spin waves or magnons are low-lying excitations and therefore $c(T)\propto T^{3/2}$ for all ferromagnets. An order-disorder phase transition occurs at the Curie temperature $T_{\rm C}$, see Table 1. Note that $T_{\rm C}$ strongly depends on the lattice geometry. Around $T_{\rm C}$, one expects the critical behavior, $c(T)\propto|T-T_{\rm C}|^{-\alpha}$, $\alpha\approx-0.12$ (3D Heisenberg universality class, see Ref. [29]). Here, the critical exponent $\alpha<0$ and the specific heat is finite at $T_{\rm C}$. In paramagnetic phase, as the temperature tends to infinity, all quantum Heisenberg feromagnets become the Curie paramagnet. Note a peculiar behavior for the hyperkagome-lattice case: $c(T)$ exhibits a sloped shoulder in the paramagnetic phase (green curves in Fig. 2). Meanwhile we do not have a rigorous and convincing explanation for that and this issue needs further study.
III.2 Antiferromagnets
III.2.1 Bipartite lattices
Figure 3: Specific heat $c(T)$ for the $S=1/2$ Heisenberg antiferromagnet on two bipartite three-dimensional lattices. Quantum Monte Carlo simulations for the number of unit cells ${\cal L}=40$ (simple cubic) and ${\cal L}=30$ (diamond) and a small symmetry breaking field $h=10^{-4}$. Inset presents $c$ as a function of $T/T_{\rm N}$.
@@ III.2.1 Bipartite lattices
Our findings for bipartite-lattice antiferromagnets are collected in the right column of Table 1 and in Fig. 3. Again, low-lying excitations are spin waves with linear dispersion relation, leading to a power-law decay of the specific heat as $T\to 0$ with the exponent $3$. With temperature increase, the antiferromagnetic order is gradually suppressed by thermal fluctuations and is completely destroyed at the Néel temperature $T_{\rm N}$ (see Table 1) and above it.
III.2.2 Frustrated lattices
Figure 4: Specific heat $c(T)$ for the $S=1/2$ Heisenberg antiferromagnet on two frustrated three-dimensional lattices. Entropy method results. Top: Pyrochlore lattice. Solid curves correspond to $e_{0}=-0.49$ [21] whereas dashed curves correspond to $e_{0}=-0.477$ [22] with $[8,9](e)$ approximant; dash-dotted curve is taken from Ref. [13]. In addition, we present numerical-linked-cluster-expansion data taken from [30] (circles) as well as simple Padé approximants $[8,9](T)$ and $[8,8](T)$ which constitute the boundary of the gray area. Bottom: Hyperkagome lattice. The ground-state energy is determined within the entropy method and we took $[10,10](e)$ approximant to obtain all these entropy-method curves; dash-dotted curves are taken from the previous study [20]. In addition, we present simple Padé approximants $[8,9](T)$ and $[10,10](T)$ which constitute the boundary of the gray area.
@@ III.2.2 Frustrated lattices
Frustrated-lattice antiferromagnets are less understood. As said above, there is no universal method to get reliable results for $c(T)$: Quantum Monte Carlo fails because of infamous sign problem, exact diagonalization or finite-temperature Lanczos methods are restricted to too small lattices in three spatial dimensions, the rotation-invariant Green’s function method, although is, in principle, applicable, contains uncontrolled approximation. High-temperature expansion series approach may be useful, but only if its applicability is extended to sufficiently low temperatures. Therefore, we have performed the entropy-method calculations, see Sec. II.2.2. Our findings are reported in Fig. 4.
@@
First of all we note that we have extended previous calculations [13, 20] by (i) including more terms (17 instead of 13 (pyrochlore) or 20 instead of 16 (hyperkagome)) and (ii) using the present value of the ground-state energy $e_{0}$ for the pyrochlore-lattice case. For the hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnet we obtain within the entropy method a more precise estimate for the ground-state energy $e_{0}\in[-0.439\,6,-0.435\,8]$ (according to Ref. [20] $e_{0}\in[-0.441,-0.435]$).
@@
For geometrically frustrated Heisenberg magnets, no magnetic order occurs until zero temperature. (Note, however, that some symmetries for the pyrochlore-lattice case are broken in the ground state, see Refs. [21, 22, 31].) From Fig. 4, the lack of a clear magnetic ordering feature in $c(T)$ is evident: $c(T)$ shows a much broader peak (bump or rounded maxima), than usually seen peak (sharp cusp) for a magnetic transition, at $T$ somewhat below $J=1$, compare Fig. 4 to Fig. 3. (Note also that the scale for $c$ is now about 5 times smaller.) Besides the main maximum of the order of $J$, there is a low-temperature maximum, to be discussed below.
@@
We have to comment what is new in Fig. 4 in comparison to Refs. [13, 20]. Concerning the pyrochlore case, in the upper panel of Fig. 4 we show the “old” entropy-method curve $c(T)$ obtained in Ref. [13] (dash-dotted, $n=13$, $e_{0}=-0.52$, $\alpha=2$, $[6,6](e)$), the result of the numerical linked cluster expansion [30] (empty circles, above $T\approx 0.25$), as well as two Padé approximants $[8,9](T)$ and $[8,8](T)$ (gray shaded area, above $T=0.35$). “New” entropy-method curves are based on $n=17$ terms in Eq. (6), $e_{0}=-0.49$ ([21], solid) or $e_{0}=-0.477$ ([22], dashed), $[8,9](e)$ approximant, and different assumptions about a law for vanishing $c(T)$ as $T\to 0$, see also Appendix A. Obvious ambiguity in the shape of $c(T)$ is related to lack of knowledge about the ground state and low-lying excitations. Concerning the hyperkagome case, in the lower panel of Fig. 4 we show, besides the entropy-method curves, two Padé approximants $[8,9](T)$ and $[10,10](T)$ (gray shaded area, above $T=0.2$). The entropy-method curves which follow from $[10,10](e)$ approximant for an assumed scenario of low-temperature behavior of $c(T)$ are accompanied by shadow areas. The boundaries of this area correspond to two (slightly) different values of $e_{0}$ which give the same (large) number of almost coinciding curves $c(T)$ obtained from Padé approximants $[u,d](e)$, for details see Ref. [20] and also Appendix B. The new curves in the lower panel of Fig. 4 (solid) are quite close to those reported in Ref. [20] (dash-dotted).
III.2.3 Low-temperature peak and a hidden energy scale
@@
Recently, a theory of a low-temperature peak of the specific heat for frustrated quantum Heisenberg antiferromagnets has been suggested in Refs. [6, 7]. These authors considered various geometrically frustrated lattices like kagome, triangular, hyperkagome, or spinel to emphasize the existence of two temperature scales which manifest themselves in two distinct, well-distinguishable peaks in $c(T)$ and argued that the low-temperature peak comes from nonmagnetic excitations similar to spin exchanges in chains of spins, whose energy is well separated from the energy of other magnetic excitations, similar to spin flips, which give rise to the high-temperature peak in $c(T)$. Both panels of Fig. 4 with the entropy-method findings for $c(T)$ agree qualitatively with such a picture. Moreover, the authors of Refs. [7, 6] also predicted the entropy of the excitations contributing to the low-temperature peak: Because the excitations that give rise to that peak are adiabatically connected to the ground states of the Ising model on the same lattice, the associated entropy $s(T^{*})$ should match the entropy of the ground states of that Ising model ${\sf s}(0)$. Here, the entropy related to the low-temperature peak may be estimated as $s(T^{*})=\int_{0}^{T^{*}}{\rm d}{\rm T}c({\rm T})/{\rm T}$, i.e., assuming that the low-temperature peak spreads up to $T^{*}$, whereas the residual ground-state entropy of the Ising model ${\sf s}(0)=\lim_{T\to 0}{\sf s}(T)$, ${\sf s}(T)=\int_{0}^{T}{\rm d}{\rm T}{\sf c}({\rm T})/{\rm T}=\ln 2-\int_{T}^{\infty}{\rm d}{\rm T}{\sf c}({\rm T})/{\rm T}$.
@@
To check this theory, we have to compare the low-temperature peak entropy for the Heisenberg model $s(T^{*})$ with the residual ground-state entropy for the Ising model ${\sf s}(0)$. The latter quantity, ${\sf s}(0)$, follows from rigorous analytical calculations in the two-dimensional case, e.g., Refs. [4, 5, 32] or from approximate analytical and numerical calculations in the three-dimensional case, e.g., Refs. [33, 34]. The former quantity, $s(T^{*})$, is related to $c(T)$, which, in turn, is known only from numerics and mainly in the two-dimensional case, e.g., Refs. [35, 19]. (In Ref. [6] the authors compute numerically $c(T)$ of a kagome cluster of $N=18$ sites.)
@@
With the obtained temperature profiles shown in Fig. 4 we may further discuss the theory of hidden energy scale of Ref. [6] using as a testing bad two three-dimensional systems. According to Fig. 4, for the entropy under the lower temperature peak we get:
$s(T^{*})\in[0.325,0.350]$ (8)
@@
Figure 5: Towards the residual ground-state entropy of the uniaxial Ising model on the pyrochlore and hyperkagome lattices. The entropy is obtained by integration from the Monte Carlo data for the specific heat (up to ${\cal L}=30$), shown in the inset. Again, $h=10^{-4}$.
@@
On the other hand, we need the residual ground-state entropy of the Ising antiferromagnet on the pyrochlore and hyperkagome lattices. To this end, we consider the $S=1/2$ models with the uniaxial Ising-like anisotropy and perform classical Monte Carlo simulations [8, 9] for the specific heat ${\sf C}(T)=\partial{\sf E}/\partial T$ to obtain the entropy ${\sf s}(T)$ per site through the formula ${\sf s}(T)=\ln 2-\int_{T}^{\infty}{\rm d}{\rm T}{\sf c}({\rm T})/{\rm T}$. Our findings are reported in Fig. 5. We obtain ${\sf s}(0)\approx 0.1$ (pyrochlore) and ${\sf s}(0)\approx 0.5$ (hyperkagome). We may also mention here Ref. [33] (see also [36]), where some calculations using Pauling’s technique to estimate the zero-point entropy were reported. Moreover, similar classical Monte simulations were presented in Ref. [34], however, for the $S=1/2$ models with the Ising-like (i.e., easy-axis) anisotropy along the direction into the center of the tetrahedron (pyrochlore) or along line joining the triangular centers (hyperkagome). The result for the hyperkagome case only very slightly exceeds the one in Fig. 5, but the result for the pyrochlore case is noticeably larger than the one in Fig. 5 and the residual ground-state entropies differ approximately by a factor of about 2. It is worth noting, that from the physical point of view the Ising-like anisotropy (spin space) is expected to agree with the crystal symmetry (real space). The uniaxial anisotropy requires a unique crystalline axis, which is not available in the three-dimensional pyrochlore and hyperkagome lattices. In contrast, the mentioned above other Ising-like anisotropies are consistent with the lattice symmetries. For more discussion on global-axis and local-axis Ising models see Ref. [37].
@@
Comparing both data for the pyrochlore-lattice case, $0.325\ldots 0.350$ versus $0.1/0.2$, and the hyperkagome-lattice case, $0.244\ldots 0.257$ versus $0.5$, we observe that these numbers are not very close. Of course, we have to note an ambiguity in estimation of $s(T^{*})$ in Eqs. (8) and (9) simply by taking $T^{*}=0.2$ and $T^{*}=0.1$. Besides, the residual ground-state entropy of the Ising models obtained by classical Monte Carlo simulations needs better accuracy. One may also ask which two-state Ising model, uniaxial or which is consistent with crystalline axis, should be examined: Both are acceptable for adiabatic connection with the Heisenberg model but have different residual ground-state entropies. Probably a most important comment comes after taking a closer look at the hidden energy scale theory of Refs. [7, 6]. The authors considered, as a starting point, the Ising model on the lattice at hand which has the degeneracy of ground states that scales exponentially with lattice size. Then, switching on the transverse-coupling Hamiltonian, which is controlled by the anisotropy parameter $\delta$, $0\leq\delta\leq 1$, lifts the ground-state degeneracy and results in a set of low-lying excitations which is well separated from the spin-flip-like excitations until $\delta$ is close to 0. These low-lying (i.e., low-energy) excitations are responsible for a low-temperature peak of $c(T)$, which is well separated from the main high-temperature peak of $c(T)$. And the entropy associated with the low-temperature peak of $c(T)$ should match the residual ground-state entropy of the Ising model on the lattice under consideration. However, it is not clear a priori whether the low-lying excitations constitute a set of states well separated from other excitations if $\delta\to 1$. To end up, a better understanding of the hidden energy scale, in particular, on the quantitative level awaits further research.
IV Discussion and summary

### v3

@@ 3.1 Ferromagnets
Figure 2: (Colour online) Specific heat $c(T)$ for the $S=1/2$ Heisenberg ferromagnet on four three-dimensional lattices. Quantum Monte Carlo simulations for the number of unit cells ${\cal L}=40$ (simple cubic) and ${\cal L}=30$ (pyrochlore, diamond, hyperkagome) and a small symmetry breaking field $h=10^{-4}$. Inset presents $c$ as a function of $T/T_{\rm C}$.
Figure 3: (Colour online) Towards a sloped shoulder in the paramagnetic phase for the hyperkagome-lattice $S=1/2$ Heisenberg ferromagnet: high-order Padé approximants $[u,d](T)$, $u+d\leqslant 20$, to the specific heat series $c(\beta)$ versus quantum Monte Carlo simulations (which were reported in figure 2).
3.2 Antiferromagnets
3.2.1 Bipartite lattices
@@
Our findings for bipartite-lattice antiferromagnets are collected in the right-hand column of table 1 and in figure 4. Again, low-lying excitations are spin waves with linear dispersion relation, leading to a power-law decay of the specific heat as $T\to 0$ with the exponent $3$. With temperature increase, the antiferromagnetic order is gradually suppressed by thermal fluctuations and is completely destroyed at the Néel temperature $T_{\rm N}$ (see table 1) and above it. Note that $T_{\rm C}<T_{\rm N}$ as expected for the quantum $S=1/2$ Heisenberg magnets on bipartite lattices, see, e.g., reference [33].
Figure 4: (Colour online) Specific heat $c(T)$ for the $S=1/2$ Heisenberg antiferromagnet on two bipartite three-dimensional lattices. Quantum Monte Carlo simulations for the number of unit cells ${\cal L}=40$ (simple cubic) and ${\cal L}=30$ (diamond) and a small symmetry breaking field $h=10^{-4}$. Inset presents $c$ as a function of $T/T_{\rm N}$.
3.2.2 Frustrated lattices
@@
Frustrated-lattice antiferromagnets are less understood. As said above, there is no universal method to get reliable results for $c(T)$: quantum Monte Carlo fails because of infamous sign problem, exact diagonalization or finite-temperature Lanczos methods are restricted to too small lattices in three spatial dimensions, the rotation-invariant Green’s function method, although is principally applicable, contains uncontrolled approximation. High-temperature expansion series approach may be useful, but only if its applicability is extended to sufficiently low temperatures. Therefore, we have performed the entropy-method calculations, see section 2.2.2. Our findings are reported in figure 5.
@@
First of all we note that we have extended previous calculations [15, 22] by (i) including more terms (17 instead of 13 (pyrochlore) or 20 instead of 16 (hyperkagome)) and (ii) using the present value of the ground-state energy $e_{0}$ for the pyrochlore-lattice case. For the hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnet we obtain within the entropy method a more precise estimate for the ground-state energy $e_{0}\in[-0.4396,-0.4358]$ (according to reference [22] $e_{0}\in[-0.441,-0.435]$).
@@
For geometrically frustrated Heisenberg magnets, no magnetic order occurs until zero temperature. (Note, however, that some symmetries for the pyrochlore-lattice case are broken in the ground state, see references [23, 24, 34].) From figure 5, the lack of a clear magnetic ordering feature in $c(T)$ is evident: $c(T)$ shows a much broader peak (bump or rounded maxima), than the usually seen peak (sharp cusp) for a magnetic transition, at $T$ somewhat below $J=1$, compare figure 5 to figure 4. (Note also that the scale for $c$ is now about 5 times smaller.) Besides the main maximum of the order of $J$, there is a low-temperature maximum, to be discussed below.
@@
We have to comment what is new in figure 5 in comparison to references [15, 22]. Concerning the pyrochlore case, in the upper panel of figure 5 we show the ‘‘old’’ entropy-method curve $c(T)$ obtained in reference [15] [dash-dotted, $n=13$, $e_{0}=-0.52$, $\alpha=2$, $[6,6](e)$], the result of the numerical linked cluster expansion [35] (empty circles, above $T\approx 0.25$), as well as two Padé approximants $[8,9](T)$ and $[8,8](T)$ (gray shaded area, above $T=0.35$). ‘‘New’’ entropy-method curves are based on $n=17$ terms in equation (2.1), $e_{0}=-0.49$ ([23], solid) or $e_{0}=-0.477$ ([24], dashed), $[8,9](e)$ approximant, and different assumptions about a law for vanishing $c(T)$ as $T\to 0$, see also appendix A. Obvious ambiguity in the shape of $c(T)$ is related to the lack of knowledge about the ground state and low-lying excitations. Concerning the hyperkagome case, in the lower panel of figure 5 we show, besides the entropy-method curves, two Padé approximants $[8,9](T)$ and $[10,10](T)$ (gray shaded area, above $T=0.2$). The entropy-method curves which follow from $[10,10](e)$ approximant for an assumed scenario of low-temperature behavior of $c(T)$ are accompanied by shadow areas. The boundaries of this area correspond to two (slightly) different values of $e_{0}$ which give the same (large) number of almost coinciding curves $c(T)$ obtained from Padé approximants $[u,d](e)$, for details see reference [22] and also appendix B. The new curves in the lower panel of figure 5 (solid) are quite close to those reported in reference [22] (dash-dotted).
Figure 5: (Colour online) Specific heat $c(T)$ for the $S=1/2$ Heisenberg antiferromagnet on two frustrated three-dimensional lattices. Entropy method results. Top: pyrochlore lattice. Solid curves correspond to $e_{0}=-0.49$ [23] whereas dashed curves correspond to $e_{0}=-0.477$ [24] with $[8,9](e)$ approximant; dash-dotted curve is taken from reference [15]. In addition, we present numerical-linked-cluster-expansion data taken from [35] (circles) as well as simple Padé approximants $[8,9](T)$ and $[8,8](T)$ which constitute the boundary of the gray area. Bottom: hyperkagome lattice. The ground-state energy is determined within the entropy method and we took $[10,10](e)$ approximant to obtain all these entropy-method curves; dash-dotted curves are taken from the previous study [22]. In addition, we present simple Padé approximants $[8,9](T)$ and $[10,10](T)$ which constitute the boundary of the gray area.
3.2.3 Low-temperature peak and a hidden energy scale
@@
Recently, a theory of a low-temperature peak of the specific heat for frustrated quantum Heisenberg antiferromagnets was suggested in references [8, 9]. These authors considered various geometrically frustrated lattices such as kagome, triangular, hyperkagome, or spinel to emphasize the existence of two temperature scales which manifest themselves in two distinct, well-distinguishable peaks in $c(T)$ and argued that the low-temperature peak comes from nonmagnetic excitations, similar to spin exchanges in the chains of spins, whose energy is well separated from the energy of other magnetic excitations, similar to spin flips, which give rise to the high-temperature peak in $c(T)$. Both panels of figure 5 with the entropy-method findings for $c(T)$ qualitatively agree with such a picture. Moreover, the authors of references [9, 8] also predicted the entropy of the excitations contributing to the low-temperature peak: since the excitations that give rise to that peak are adiabatically connected to the ground states of the Ising model on the same lattice, the associated entropy $s(T^{*})$ should match the entropy of the ground states of that Ising model s(0){s}(0). Here, the entropy related to the low-temperature peak may be estimated as $s(T^{*})=\int_{0}^{T^{*}}{\rm d}Tc(T)/T$, i.e., assuming that the low-temperature peak spreads up to $T^{*}$, whereas the residual ground-state entropy of the Ising model ${s}(0)=\lim_{T\to 0}{s}(T)$, ${s}(T)=\int_{0}^{T}{\rm d}T{c}(T)/T=\ln 2-\int_{T}^{\infty}{\rm d}T{c}(T)/T$.
@@
To check this theory, we have to compare the low-temperature peak entropy for the Heisenberg model $s(T^{*})$ with the residual ground-state entropy for the Ising model s(0){s}(0). The latter quantity, s(0){s}(0), follows from rigorous analytical calculations in the two-dimensional case, e.g., references [6, 7, 36] or from approximate analytical and numerical calculations in the three-dimensional case, e.g., references [37, 38]. The former quantity, $s(T^{*})$, is related to $c(T)$, which, in turn, is known only from numerics and mainly in the two-dimensional case, e.g., references [39, 21]. (In reference [8] the authors numerically compute $c(T)$ of a kagome cluster of $N=18$ sites.)
@@
With the obtained temperature profiles shown in figure 5 we may further discuss the theory of the hidden energy scale of reference [8] using as a testing bed two three-dimensional systems. According to figure 5, for the entropy under the lower temperature peak we get:
$s(T^{*})\in[0.325,0.350]$ (3.1)
@@
On the other hand, we need the residual ground-state entropy of the Ising antiferromagnet on the pyrochlore and hyperkagome lattices. To this end, we consider the $S=1/2$ models with the uniaxial Ising-like anisotropy and perform classical Monte Carlo simulations [10, 11] for the specific heat ${C}(T)=\partial{E}/\partial T$ to obtain the entropy s(T){s}(T) per site through the formula ${s}(T)=\ln 2-\int_{T}^{\infty}{\rm d}T{c}(T)/T$. Our findings are reported in figure 6. We obtain ${s}(0)\approx 0.1$ (pyrochlore) and ${s}(0)\approx 0.5$ (hyperkagome). We may also mention here reference [37] (see also [40]), where some calculations using Pauling’s technique to estimate the zero-point entropy were reported. Moreover, similar classical Monte Carlo simulations were presented in reference [38], although for the $S=1/2$ models with the Ising-like (i.e., easy-axis) anisotropy along the direction into the center of the tetrahedron (pyrochlore) or along the line joining the triangular centers (hyperkagome). The result for the hyperkagome case just very slightly exceeds the one in figure 6, but the result for the pyrochlore case is noticeably larger than the one in figure 6 and the residual ground-state entropies differ approximately by a factor of about 2. It is worth noting that from the physical point of view the Ising-like anisotropy (spin space) is expected to agree with the crystal symmetry (real space). The uniaxial anisotropy requires a unique crystalline axis, which is not available in the three-dimensional pyrochlore and hyperkagome lattices. By contrast, the other above mentioned Ising-like anisotropies are consistent with the lattice symmetries. For more discussion on global-axis and local-axis Ising models see reference [41].
Figure 6: (Colour online) Towards the residual ground-state entropy of the uniaxial Ising model on the pyrochlore and hyperkagome lattices. The entropy is obtained by integration from the Monte Carlo data for the specific heat (up to ${\cal L}=30$), shown in the inset. Again, $h=10^{-4}$.
@@
Comparing both the data for the pyrochlore-lattice case, $0.325\ldots 0.350$ versus $0.1/0.2$, and the hyperkagome-lattice case, $0.244\ldots 0.257$ versus $0.5$, we observe that these numbers are not very close. Of course, we have to note an ambiguity in estimation of $s(T^{*})$ in equations (3.1) and (3.2) simply by taking $T^{*}=0.2$ and $T^{*}=0.1$. Besides, the residual ground-state entropy of the Ising models obtained by classical Monte Carlo simulations needs a better accuracy. One may also ask which exactly two-state Ising model, uniaxial or the one consistent with crystalline axis, should be examined: Both are acceptable for adiabatic connection with the Heisenberg model but have different residual ground-state entropies. Probably the most important comment comes after taking a closer look at the hidden energy scale theory of references [9, 8]. The authors considered, as a starting point, the Ising model on the lattice at hand which has the degeneracy of ground states that scales exponentially with the lattice size. Then, switching on the transverse-coupling Hamiltonian, which is controlled by the anisotropy parameter $\delta$, $0\leqslant\delta\leqslant 1$, lifts the ground-state degeneracy and results in a set of low-lying excitations which is well separated from the spin-flip-like excitations until $\delta$ is close to 0. These low-lying (i.e., low-energy) excitations are responsible for a low-temperature peak of $c(T)$, which is well separated from the main high-temperature peak of $c(T)$. And the entropy associated with the low-temperature peak of $c(T)$ should match the residual ground-state entropy of the Ising model on the lattice under consideration. However, it is not clear a priori whether the low-lying excitations constitute a set of states well separated from other excitations if $\delta\to 1$. To end up, a better understanding of the hidden energy scale, in particular, at the quantitative level awaits further research.
4 Discussion and summary